DETECTION AND FORMATION OF INTERSTELLAR c-C_{3}D_{2}

\underline{SILVIA SPEZZANO}, SANDRA BRÜNKEN, PETER SCHILKE, I. Physikalisches Institut, Universität zu Köln, Zülpicher Str. 77, 50937 Köln, Germany; KARL M. MENTEN, Max-Planck-Institut für Radioastronomie, Auf dem Hügel 69, 53121 Bonn, Germany; PAOLA CASELLI, School of Physics and Astronomy, University of Leeds, Leeds LS2 9JT, UK; MICHAEL C. McCARTHY, Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138, and School of Engineering & Applied Sciences, Harvard University, 29 Oxford St., Cambridge, MA 02138; LUCA BIZZOCCHI, Centro de Astronomia e Astrofísica, Observatório Astronómico de Lisboa, Tapada da Ajuda, 1349-018 Lisboa, Portugal; SANDRA TREVINO, IRAM, 18012, Granada, Spain; YURI AIKAWA, Department of Earth and Planetary Sciences, Kobe University, Kobe 657-8501, Japan; and STEPHAN SCHLEMMER, I. Physikalisches Institut, Universität zu Köln, Zülpicher Str. 77, 50937 Köln, Germany.

Multiply deuterated molecules are unique observational probes for the earliest stages of star formation. Cyclopropenylidene, c-C_{3}H_{2}, is an ideal probe for deuterium chemistry. It is one of the most widespread molecules in our Galaxy and it has the possibility of double deuteration. Furthermore, since c-C_{3}H_{2} is an early-type molecule, it is a particular useful tool to investigate early stages of a molecular cloud. This makes observations of its deuterated forms particularly important to test time-dependent chemical codes which include deuteration processes. The centimeter and millimeter wavelength spectra of doubly deuterated cyclopropenylidene have recently been measured in the laboratory^{a}, allowing for the first time a search for c-C_{3}D_{2} in space. We report the detection of c-C_{3}D_{2} in two starless cores, L1544 and TMC-1C. The deuteration of this small hydrocarbon ring is analysed with a comprehensive gas-grain model, the first including doubly deuterated species. The observed abundances of c-C_{3}D_{2} can be explained solely by gas-phase processes, supporting the idea that c-C_{3}H_{2} is a good indicator of gas-phase deuteration.

aS. Spezzano, F. Tamassia, S. Thorwirth, P. Thaddeus, C. A. Gottlieb, and M. C. McCarthy Astroph. J. Supp. Series \underline{200}(1),